# Learning Plasma Dynamics and Robust Rampdown Trajectories with Predict-First Experiments at TCV

### v1

@@ 2.2 Medium-Scale Validation of NSSM Predictions
Actions $\mathbf{a}_{t}$
$dI_{p}/dt$ Plasma current ramp-rate MA/s
$P_{NBI}$ Neutral beam injection (NBI) heating MW
$d\kappa/dt$ Elongation ramp-rate 1/s
$da_{minor}/dt$ Minor radius ramp-rate m/s
Table 1: The set of observations predicted by the learned dynamics model in response to action inputs. The time derivatives of certain quantities are chosen as the action to allow the ramp-rate to be constrained during trajectory optimization. Note that certain actions are not optimized in this work, but are important to include for the purpose of training the dynamics model; these actions are simply set to zero during the trajectory optimization process.
Figure 4: Model prediction accuracy as a function of prediction horizon during rampdowns on the validation set of 131 shots. Both individual data points and percentiles are shown. Figure 16 in the Supplementary Information shows similar model performance on the smaller scale fine tuning dataset.
Parameter Description Units
Inputs
$W_{tot}$ Plasma total stored energy $10^{4}$ J
$\bar{n}_{e,20}$ Line-averaged electron density (measured by interferometry) $10^{20}\,\text{m}^{-3}$
$P_{NBI}$ Neutral beam injection (NBI) heating MW
$P_{ECRH}$ Electron cyclotron resonance heating (ECRH) injected power MW
$I_{p}$ Plasma current $10^{5}$ A
$\kappa$ Elongation [-]
$a_{minor}$ Minor radius m
$\delta$ Triangularity [-]
Predictions
$T_{e}(\rho)$ Electron temperature profile keV
$n_{e,20}(\rho)$ Electron density profile $10^{20}\,\text{m}^{-3}$
Table 2: Input and output quantities for the profile predictor.
@@
2.3 Preventing VDEs by Designing for Robustness to Control Error
@@
Figure 6: Comparison of a shot designed prior to adding control uncertainty, #81751, and after, #82875. We see that the vertical growth rate $\gamma_{vgr}$ is highly sensitive to gap control error in #81751, but similar control errors experienced in #82875 result in negligible changes in stability. Time shown is relative to 1.4 seconds for #81751 and 1.3 seconds for #82875 to align the moment the two shots experience similar control errors.
2.4 Improving Terminations by Incremental Re-training
Both the NSSM and trajectories were incrementally re-trained on newly generated data from experimental run-days, which resulted in more robust and faster rampdowns for the baseline high performance scenario, as shown in Figure 7. The speed of the model enabled re-training of both the model and trajectories in fewer than ten hours total on a single A100 GPU. The un-optimized solution in #81101 involved a NBI power rampdown while keeping constant plasma current, to allow for a decrease in density to avoid the Greenwald limit, a solution which the RL approach initially decided on as well, as shown with #81751, with an even more conservative current ramp and introducing a reduction in $\kappa$. As discussed in 2.3, this shot resulted in a VDE, and, with the introduction of an uncertainty distribution on $g_{HFS}$, the solution in #81830 resulted in less of a reduction in the minor radius $a_{minor}$, which helped eliminate the $\gamma_{vgr}$ spikes. Subsequent dynamics model training and trajectory optimization resulted in a solution in #82876 which allows for an immediate reduction in $I_{p}$ without running into a density limit, highlighting the ability for the workflow to assist in gradually making improvements. All optimized trajectories involved a fast initial drop in $P_{NBI}$, followed by a slower ramp phase, although the rates and critical points for the transition differed from shot to shot.
Figure 7: Experimental traces of key actuators (left) and relevant quantities (right) for the baseline high performance scenario over the course of the experiment. Time is set relative to the beginning of the termination phase.
2.5 Predict-First Results for the Extrapolation Test
As shown in Figure 8, experimental results from #82878 largely fell within this distribution, with accurate predictions of the stored energy and density dynamics. Arguably the largest sources of error came from unreliable control of the plasma shape, contributing to errors in quantities like the rotational transform $\iota_{95}\equiv\frac{1}{q_{95}}$, and also leading to an earlier than expected HL back-transition. #82878 also started near the density limit, a challenging situation which motivated the introduction of a delay to the $I_{p}$ ramp in the baseline scenario, but the RL algorithm was able to determine a trajectory to immediately decrease $I_{p}$, which is desirable, while keeping $f_{GW}$ roughly constant. #82877 fell further out of distribution due to a loss of NBI power, and the presence of a neo-classical tearing mode (NTM) at the beginning of rampdown that did not exist in TCV #82878, as shown by Figure 17 in the Supplementary Information. Fortunately, these un-modeled off-normal events (ONEs) did not take the plasma far enough out of distribution to cause a disruption. As discussed earlier, the profile predictor was removed to help accelerate trajectory optimization, but post shot evaluation of the profile predictor on the 0D scalars generated by the training environment, shown in Figure 9, shows reasonable agreement against experimental Thomson measurements.
@@ 2.5 Predict-First Results for the Extrapolation Test
Figure 8: Action trajectories and a priori predictions of plasma dynamics during rampdown, compared to experimental results from TCV #82877 and #82878. The RL training environment accounts for uncertainty in actuation with distributions on the action trajectory; the average of the distribution is used for shot programming. Control of the plasma shape proved to be a challenge for this phase, an issue also observed in previous rampdown studies [9, 7].
Figure 9: Post-hoc predictions of kinetic profiles compared to Thomson Scattering measurements.
3 Discussion
Our results demonstrate that the developed approach to learning plasma dynamics can predict the highly transient rampdown phase with a modest dataset and even make small extrapolations to higher performance regimes. The relative sample efficiency of the approach, only requiring five shots in the relevant high performance regime, indicates this may be a viable approach for upcoming tokamaks like SPARC and ITER, which will initially operate at low performance before incrementally increasing performance. Developing robust terminations during such incremental campaigns is crucial, as highlighted by the 2020 JET high performance campaign where a 15% increase in plasma current, from 3MA and 3.5MA, raised disruptivity from $\approx 20\%$ to $\approx 50\%$ [12]. Prediction metrics on the validation dataset, as shown in Figure 4, shows this approach yields accurate predictions for the majority of ramp-downs, but the 5% worst cases can involve large prediction errors, meriting further investigation.
The developed RL approach for designing robust trajectories yielded promising improvements in the plasma current and stored energy at time of termination, with incremental re-training improving the ramp speed. This result represents one of the first successful demonstrations of trajectory design with robustness to real-world uncertainties for tokamaks, which has historically been infeasible due to the computational cost of simulation. A degree of statistical significance is shown, but the sample size is still relatively small; a larger scale study would more thoroughly determine the efficacy of the approach. Although a large set of uncertainties was accounted for, detailed further in Table 4, experimental results involved additional uncertainties, such as the NBI failure in #82877, that still need to be addressed to further improve the robustness of trajectories. Robustness to hardware failure is of particular interest for future work as an exhaustive survey of disruption causes at JET has revealed hardware failure as a significant contributing factor to disruptions [37]. It is also noteworthy that the RL designed action trajectories tended to be relatively simple, suggesting that the key important ingredient is the fast and parallelized simulation model, as a human expert may be able to find similar trajectories if given access to the simulation model.
4 Methods
4.1 The Neural State-Space Model
$\dot{\mathbf{x}}$ $=A\mathbf{x}+B\mathbf{a}$ (1a)
$\mathbf{o}$ $=C\mathbf{x}+D\mathbf{a}$ (1b)
$\dot{\mathbf{x}}=NN_{\theta}(\mathbf{x})$ (2)
$\dot{\mathbf{x}}(t)$ $=f_{\theta}(\mathbf{x},\mathbf{a})$ (3a)
$\mathbf{o}(t)$ $=O_{\theta}(\mathbf{x},\mathbf{a})$ (3b)
Figure 10: The training process of a NSSM, showing quantities from the experimental database, computed quantities, and functions. The loss functional is defined as the time-integrated value of an instantaneous loss function $l$, and $\nabla_{\theta}\mathcal{L}$ denotes the gradient of the parameter vector w.r.t. the value of the loss functional.
4.1.1 The Dynamics Function $f_{\theta}(\mathbf{x},\mathbf{a})$
We begin by defining the following confinement laws:
$\text{tanhHeaviside}(x)$ $\equiv\frac{1}{2}(\tanh(kx+1))$ (4e)
where the parameters to be learned include all coefficients $c_{*}$ and neural network parameters. The laws are structured to multiply a portion corresponding to L-mode, a neural network correction factor, and an H-mode correction factor. The L-mode term reflects standard confinement scalings, but with the introduction of a $\dot{I}_{p}$, which was found to help better capture short-term effects of ramping plasma current. The neural network output includes a tanhClip final activation that constrains it’s output to the range $[0.75,1.25]$, thus controlling the maximum adjustment the network is allowed to make. The hmode function includes a tanhHeaviside function which provides a smooth transition between one to zero once $P_{input}$ falls below the learned back-transition threshold, which is structured to reflect the Martin scaling [41]. Note that the use of the hmode function output as a power de-activates the H-mode correction term once hmode transitions from one to zero. While, in principle, the neural network should be able to learn the effects of H-mode implicitly, we found that adding an explicit H-mode correction factor helped improve model predictions in our low-data regime. The $k$ factor controls the smoothness of both the tanhClip and tanhHeaviside functions.
where $\sigma$ is the sigmoid function, $NN_{ohm,rad}$ is a network that predicts two quantities; the first is multiplied by $I_{p}^{2}$ to serve as an Ohmic heating term, and the second is multiplied by density and volume to serve as the radiated power term. $NN_{wall}$ is included to account for possible wall fueling effects when in a limited configuration, and is multiplied by an exponential in the HFS gap to de-activate it when diverted. Additional simple constants are included to account for fueling from both NBI and gas puffing. We note that, in both cases, the included terms do not capture important state dependencies and time delays, but they proved sufficient for this use case. All NNs used in the dynamics function $f$ are simple MLPs with GELU activations on the hidden state and tanhClip functions as final activations to constrain their outputs to reasonable ranges. The dynamics of density times volume are predicted; in cases where density itself is used (e.g. to compute the Greenwald Fraction), the following volume approximation is used to recover density:
4.1.2 The Observation Function $O_{\theta}(\mathbf{x},\mathbf{a})$
where $\beta_{p}$ is computed in accordance to the LIUQE definition [42], $f_{GW}$ is the usual Greenwald Fraction [10], $q_{95}$ is the approximation given in [43] with the squareness factor set to 1, $NN_{vgr}$ is a MLP with GELU activation and a scaled sigmoid output, and $NN_{prof}$ is a neural-operator based profile predictor, discussed further in the next subsection.
4.1.3 Neural Operator Based Profile Predictor
Figure 11: (Left) explained variance as a function of number of PCA modes, showing high explained variance with a relatively low dimensional representation of kinetic profiles on a database of 1341 shots. (Right) an example of kinetic profiles and their corresponding approximation with three PCA modes.
$f_{out}=\mathcal{F}_{\theta}(f_{in})$ (8)
And we predict these profiles using a set of 0D scalars, where every scalar is a control parameter except stored energy $W_{tot}$ and $\bar{n}_{e,20}$. The full set of input and output parameters is listed in Table 2. During training, the $\rho$ grid corresponding to the dataset is chosen to evaluate the basis functions, but arbitrary alternative grids can be used during inference time.
4.2 Training Data Distribution
The dataset used for training models in this work consists of 442 shots with rampdowns that are at least partially complete, gathered with DEFUSE (Disruption and Event analysis framework for FUSion Experiments) [46]. The initial training phase involved training on 311 shots of data, with the rest of the dataset used for validation. After the initial training phase, the model is further trained on a fine-tuning dataset of 44 shots, during this phase all of the model weights except those in the $\tau_{E}$ and $\tau_{N}$ hybrid confinement laws described in 4.1.1 are frozen. As shown in Figure 12, the dataset consists of only five shots of data anywhere near the relevant high-performance region.
Figure 12: The data distribution used for training the NSSM, visualized in ($I_{p}$, $f_{GW}$) space and ($I_{p}$, $\beta_{N}$) space.
4.3 Reward Function
One challenge with specifying constraint limits is the difference in severity of violating different constraints, and the, at times, weak correlations between physical quantities and disruptions. To address this issue, we partition constraints into “soft” constraints, which incur a small penalty to discourage, but not forbid, the algorithm from finding solutions that violates these limits, and “hard” constraints which incur a large penalty to strictly enforce constraint violation. We note that while methods in the constrained optimization literature often mathematically express constraints separately from the objective function being optimized, most practical implementations of constrained optimization algorithms enforce constraints by rewriting constraints as penalty terms in the objective function [48, 49], an approach we also adopt. Stochastic optimization across a distribution of outcomes introduces a challenge: trying to avoid limits for every scenario will likely result in excessively slow and conservative solutions [50], which itself poses it’s own risk. To address this challenge, we utilize chance-constraints, a technique often utilized in the autonomous driving literature [51, 52], and only activate the constraint if violation probability exceeds a certain threshold. In this set of experiments, this threshold is chosen as 5%. Reward function parameters used for the final four shots are shown in Table 3.
Category Parameter Value
Hard Limits $f_{GW}$ 1.0
$\iota_{95}$ 0.5
Soft Limits $f_{GW}$ 0.8
$\beta_{p}$ 1.75
$\gamma_{vgr}$ 0.75
$\iota_{95}$ 0.313
Parameters $c_{time}$ 5.0
$c_{I_{p}}$ 1.0
$c_{W}$ 1.0
$c_{soft}$ 1.0×103
$c_{hard}$ 5.0×104
Table 3: Reward function parameters. Note that the values corresponding to limits are the maximum allowed values.
4.4 Uncertainty Model
In experimental reality, the time evolution of plasma dynamics is highly nonlinear and subject to considerable amounts of uncertainty, as evidenced by the two same-scenario shots shown in Figure 8 which begin at drastically different initial conditions. To design trajectories that have robustness to large variability and off-normal events, we defined an uncertainty model for the RL training environments, and sampled from this uncertainty model for each training environment used during training. The uncertainty model includes random variables for both the initial state of the plasma during rampdown and disturbances/model uncertainties that affect the time-varying plasma dynamics. To account for the fact that accidental H-L back-transition implies the initial state of the plasma may start in either H or L-mode, the initial state distribution is modeled as a bi-modal mixture model, with a 50% chance of any given RL training environment starting in either H or L-mode. In some cases, uncertainty distributions could easily be quantified from past experimental data (such as tracking error in the plasma current), or from model prediction accuracy (such as $\gamma_{vgr}$), but in other cases the distribution was chosen in ad-hoc fashion, upon identifying additional sources of uncertainty in experiment. Table 4 summarizes the random variables, parameterized distributions, and quantification methods used in this work. As discussed in Section 2.5, this uncertainty model proved to be non-exhaustive in experiment, further highlighting the need to advance experimental uncertainty quantification and robust control in the context of fusion plasma control.
@@ 4.4 Uncertainty Model
Table 4: The uncertainty model used in RL training environments to design trajectories with robustness to distributional uncertainty.
4.5 RL Methods
$\mathbf{a}=\pi(\mathbf{o})$ (11)
$\mathbf{a}=\pi(t)$ (12)
After an initial trial with Proximal Policy Optimization (PPO) [53], we found OpenAI-ES, an evolutionary strategy (ES) designed for policy optimization, to work better in practice [54]. This is possibly explained by the theoretical analysis given in the paper introducing OpenAI-ES, which suggests that RL problems with long time horizons and actions that have long-lasting effects may be better solved with ES approaches than the dominant paradigm of policy gradient methods [54].
4.6 Deployment to TCV
5 Acknowledgments
This work was funded in part by Commonwealth Fusion Systems. This work has been carried out within the frame-work of the EUROfusion Consortium, via the Euratom Research and Training Programme (Grant Agreement No 101052200 - EUROfusion) and funded by the Swiss State Secretariat for Education, Research, and Innovation (SERI). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission, or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them. The authors would like to acknowledge the Engaging cluster, managed by the MIT Office of Research Computing and Data, which was used in this work for model training.
6 Author Contributions
Allen M. Wang led the project, developed the dynamics modeling approach and RL problem formulation, and led writing of the paper. Alessandro Pau developed the machine learning dataset, high performance integrated control scenario, and integration of the approach into TCV. Oswin So and Charles Dawson worked with Allen M. Wang to develop the RL training environments and methods. Olivier Sauter enabled TCV integration, and identified and debugged the radial control issue. Cristina Rea managed the collaboration and advised the project on disruptions and machine learning, contributed to manuscript drafting and revisions. Mark D. Boyer advised the project from a controls perspective and motivated the approach taken for dynamics modeling. Anna Vu and Cristian Galperti enabled integration with the integrated control system. Chuchu Fan advised Allen M. Wang, Oswin So, and Charles Dawson on controls and RL approaches. Antoine Merle helped with MHD equilibrium codes. Yoeri Poels and Cristina Venturini helped with the TCV dataset. Stefano Marchioni developed the growth rate calculation method used in this work.
7 Data Availability
Data to generate figures found in this paper are available in the source code repository. The complete dataset used for training and validation can be obtained by contacting Alessandro Pau or Allen M. Wang.
8 Code Availability
Source code is provided in the supplementary materials.

### v2

@@ Acknowledgments
Actions $\mathbf{a}_{t}$
$dI_{p}/dt$ Plasma current ramp-rate MA/s
$P_{NBI}$ Neutral beam injection (NBI) heating MW
$d\kappa/dt$ Elongation ramp-rate 1/s
$da_{minor}/dt$ Minor radius ramp-rate m/s
@@
Table 1: The set of observations predicted by the learned dynamics model in response to action inputs. The time derivatives of certain quantities are chosen as the action to allow the ramp-rate to be constrained during trajectory optimization. Note that certain actions are not optimized in this work, but are important to include for the purpose of training the dynamics model; these actions are simply set to zero during the trajectory optimization process. Also note that the set of inputs used for the profile predictor component is $[W_{tot},\bar{n}_{e,20},P_{NBI},P_{ECRH},I_{p},\kappa,a_{minor},\delta]$.
Figures
Figure 1: Examples of plasma disruptions during rampdowns, and an example non-disruptive result obtained after deployment of the developed method. a) Data from an illustrative TCV pulse (#62370), showing the rampup, flattop, and rampdown phases, which are defined by the plasma current $I_{p}$. The bottom subplot shows how the rampdown pushes the plasma closer to an instability limit, in this case the Greenwald density limit, defined by the Greenwald fraction $f_{GW}=1$. Note this limit is approximate due to the, at present, incomplete physics understanding of the density limit [73, 74]. Also note that the flattop phase is abbreviated here to more clearly highlight the rampup and rampdown. b) A comparison of the current, $I_{p}$, stored energy, $W_{tot}$, and neutral beam injection power, $P_{NBI}$, trajectories for a baseline shot and an optimized shot, showing a faster, and non-disruptive, decrease in the plasma current and stored energy relative to the baseline.
Figure 2: Overview of the methodology and key statistical results. a) The training data distribution, with a modest dataset size at low performance and very few shots in the relevant high performance regime. The target scenarios for this work at 140kA and 170kA high normalized performance are shown. b) Depiction of the dynamics model training method, which involves comparing results from forward simulation of a NSSM against experimental data to compute the gradient of loss with respect to model parameters. c) Depiction of the trajectory optimization process. In addition to the trained dynamics model, the reinforcement learning (RL) training environment is defined by a reward function specifying the desired goal and a set of random variables that training environments are parallelized against, to find a trajectory that has robustness to uncertainties and off-normal events. d) Scatter plot of plasma current, $I_{p}$, and stored energy, $W_{tot}$, at time of plasma termination. Bottom-right table shows p-values from the Mann-Whitney U test comparing performance of experimental shots, with and without debug shots included, relative to the control set of all shots in the database with $\beta_{N}$ > 1.5.
Figure 3: A shot by shot breakdown of every shot in the experiment. The results show the plasma current, $I_{p}$, and stored energy, $W_{tot}$, at time of plasma termination along with additional contextual information. The arbitrary shading is for distinguishing groups of shots described by the labels.
Figure 4: Model validation metrics. Model prediction accuracy as a function of prediction horizon during rampdowns on the validation set of 131 shots. Both individual data points and percentiles are shown. Shown quantities are the line-averaged electron density times volume, $\bar{n}_{e,20}V$, the total stored energy, $W_{tot}$, the Greenwald Fraction, $f_{GW}$, the poloidal beta, $\beta_{p}$, the vertical instability growth rate, $\gamma_{vgr}$, and the safety factor at the 95th flux surface, $q_{95}$. Figure S4 in the Supplementary Information shows similar model performance on the smaller scale fine tuning dataset.
@@
Figure 6: Demonstration of increased robustness to control uncertainty. a) shows a comparison of a shot designed prior to adding control uncertainty, #81751, and after, #82875. We see that the vertical instability growth rate, $\gamma_{vgr}$, is highly sensitive to control error in the high field side gap, $g_{HFS,error}$, in #81751, but similar control errors experienced in #82875 result in negligible changes in stability. Time shown, $t_{relative}$, is the time relative to 1.4 seconds for #81751 and 1.3 seconds for #82875 to align the moment the two shots experience similar control errors. Additional quantities shown are the plasma elongation $\kappa$, minor radius $a_{minor}$, high-field side gap, $g_{HFS}$, and low-field side gap, $g_{LFS}$. b) shows $\gamma_{vgr}$ distributions under 1.5cm of minor radius control error for #81751 and #82875. Control error was simulated by introducing variations in the control points provided to the free-boundary equilibrium solver, FBT [70, 75], used for shot preparation. The resulting equilibria were then input into the $\gamma_{vgr}$ computation method used in this work [41].
Figure 7: Incremental improvements for the baseline high performance scenario. Experimental traces of key actuators are shown (left), including plasma current, $I_{p}$, neutral beam power inject, $P_{NBI}$, elongation, $\kappa$, and minor radius, $a_{minor}$. Additional relevant quantities (right) include the stored energy, $W_{tot}$, line-averaged electron density, $\bar{n}_{e,20}$, Greenwald fraction, $f_{GW}$, and vertical instability growth rate, $\gamma_{vgr}$. Time is set relative to the beginning of the termination phase. The plasma current, $I_{p}$, and stored energy, $W_{tot}$, trajectories show improvement in rampdown speed and disruptivity as the experiment progressed.
@@
Figure 8: A priori predictions and experimental results for the extrapolation test scenario. a) Action trajectories and a priori predictions of plasma dynamics during rampdown, compared to experimental results from TCV #82877 and #82878. The RL training environment accounts for uncertainty in actuation with distributions on the action trajectory; the average of the distribution (in solid black) is used for shot programming. Control of the plasma shape proved to be a challenge for this phase, an issue also observed in previous rampdown studies [9, 7]. Shown action variables include the plasma current, $I_{p}$, neutral beam injected power, $P_{NBI}$, minor radius, $a_{minor}$, elongation, $\kappa$, and high-field side gap, $g_{HFS}$. Shown predictions and constraints include the stored energy, $W_{tot}$, the line-averaged electron density, $\bar{n}_{e,20}$, poloidal beta, $\beta_{p}$, Greenwald fraction, $f_{GW}$, rotational transform at the 95% flux surface, $\iota_{95}$, and vertical instability growth rate, $\gamma_{vgr}$. b) Post-hoc predictions of electron temperature, $T_{e}$, and density, $n_{e}$, profiles compared to Thomson Scattering measurements.
Figure 9: Depiction of the Neural State-Space Model (NSSM) training methodology. The NSSM, defined by the dynamics function $f_{\theta}$ and observation function $O_{\theta}$ with parameters $\theta$, is simulated forward in time, given an initial state $\mathbf{x}_{0}$ and an action trajectory $\mathbf{a}_{0:T}$, to generate a sequence of simulated observations, $\hat{\mathbf{o}}_{0:T}$. The simulated observations are compared with experimental observations via the loss functional $\mathcal{L}$, which is defined as the time-integrated value of an instantaneous loss function $l$. Adjoint methods in diffrax [18] and JAX automatic differentiation then yields the gradient of model parameters with respect to loss, $\nabla_{\theta}\mathcal{L}$, which allows the optimizer to update the parameters $\theta$.
1 Supplementary Information
@@
Table S2: RL uncertainty model. The uncertainty model used in RL training environments to design trajectories with robustness to distributional uncertainty for both the 140kA and 170kA scenarios.
Category Parameter Value
Optimizer Initial Learning Rate 0.02
Final Learning Rate 0.0001
Transition Steps 200
Decay Rate 0.9
Weight Decay 0.005
$NN_{ohm,rad}$ Depth 2
Width 32
$NN_{vgr}$ Depth 3
Width 256
$NN_{conf}$ Depth 1
Width 128
$NN_{prof}$ Depth 2
Width 64
Category Parameter Value
Hard Limits $f_{GW}$ 1.0
$\iota_{95}$ 0.5
Soft Limits $f_{GW}$ 0.8
$\beta_{p}$ 1.75
$\gamma_{vgr}$ 0.75
$\iota_{95}$ 0.313
Parameters $c_{time}$ 5.0
$c_{I_{p}}$ 1.0
$c_{W}$ 1.0
$c_{soft}$ 1.0×103
$c_{hard}$ 5.0×104
Table S3: Hyperparameters and reward function parameters. (Left) Hyperparameters for the final trained model. (Right) Reward function parameters. Note that the values corresponding to limits are the maximum allowed values.
Figure S1: Debugging the legacy radial observer issue. Plot of planned radial magnetic position $r_{mag}$ against observed values from equilibrium reconstruction with LIUQE, showing exceptionally poor control performance during initial rampdown attempts. The issue was identified as a legacy $\frac{1}{I_{p}^{2}}$ term used in generating the radial position observer feed-forward trajectory, with the issue finally fixed in TCV#81751.
Figure S2: Results from the 140kA Scenario. Shown are a priori predictions of the RL environment and experimental outcomes of the final two 140kA shots.
Figure S3: Absolute errors for $\gamma_{vgr}$. Errors are shown for the initial validation set of 131 shots, showing relatively low absolute prediction error in the majority of shots.
Figure S4: Fine-tuning dataset metrics. Model prediction accuracy as a function of prediction horizon during rampdowns on the fine-tuning dataset of 44 shots.
Figure S5: NTM driven discrepancy between #82877 and #82878. Shown is the $n=1$ root mean squared (RMS) signal derived from magnetic diagnostics indicating the presence of a neo-classical tearing mode at the beginning of ramp-down for #82877 that is not present in #82878.
Figure S6: #81101 counterfactual analysis. A counterfactual analysis, using the action trajectories from the disruptive rampdown, #81101, in the RL training environment used for trajectory design. We see that the training environment predicts a large $\gamma_{vgr}$, well within the soft limit region right before the shot disrupted in experiment.
Figure S7: Principal component analysis (PCA) showing the empirical low dimensionality of kinetic profiles measured by Thomson Scattering. (Left) explained variance as a function of number of PCA modes, showing high explained variance with a relatively low dimensional representation of kinetic profiles on a database of 1341 shots. (Right) an example of kinetic profiles and their corresponding approximation with three PCA modes.
Figure S8: The data distribution used for training the NSSM. Each point corresponds to a shot, visualized in ($I_{p}$, $f_{GW}$) space and ($I_{p}$, $\beta_{N}$) space.
Figure S9: Definition of the beginning and end of the termination phase for every shot in this experiment. Note that in 81830, a low level software issue triggered an unexpected spike in plasma current at the very end of the rampdown. We define the termination time for that shot as the lowest achieved plasma current, as it is expected that net energy tokamaks where disruptions are of concern will have protection systems that prevent such incidents.